OpenStudy (anonymous):

Simplify. 9 squared over 9 squared^6 Help please? Don't tell me the answer, I just need help figuring it out (I'm terrible at math >.<)

OpenStudy (rational):

use below exponent property : \[\large \frac{\color{blue}{a}^m}{\color{blue}{a}^n}=\color{blue}{a}^{m-n}\]

OpenStudy (anonymous):

I hate to ask how, but how do I use that?

OpenStudy (rational):

\[\large \frac{\color{blue}{9}^2}{\color{blue}{9}^6}=?\]

OpenStudy (rational):

make ur best guess

OpenStudy (anonymous):

1 over 6561 ???

OpenStudy (rational):

You may be correct, but don't be so fast lets work it step by step.. first apply that exponent property

OpenStudy (anonymous):

Exponent property?

OpenStudy (rational):

\[\large \frac{\color{blue}{9}^2}{\color{blue}{9}^6}=\color{blue}{9}^{2-6}=\color{blue}{9}^{-4}\]

OpenStudy (rational):

That equals \[\large \frac{1}{\color{blue}{9}^4}\]

OpenStudy (rational):

that indeed equals 1/6561 so you're correct!

OpenStudy (anonymous):

So then how is it one of the answers?

OpenStudy (rational):

Oh could you post the screenshot of original question also